how can 56000 be invested part at 4% annual simple interest ant the remainder at 10% annual simple interest so that the interest erned by the two accounts is equal at the end if the year
Answer by mananth(16946)

x+y=56000
x@ 4%
y@10%
0.04x = 0.1y
0.04x-0.1y=0
x+y=56000 .............1
0.04x-0.1y=0 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 10
1 x 1 y = 56000
0.4 x + -1 y = 0
Add the two equations
1.4 x = 56000.000
/ 1.4
x = 40000
plug value of x in (1)
1 x + 1 y = 56000
40000 + 1 y = 56000
1 y = 56000 -40000
1 y = 16000
y = 16000
4000@4%
16000@10%